OpenStudy (anonymous):

what are the evaluate for each expression 339+638 1801+73 46.8+780 381-137 71,715-4812 930-38.7

OpenStudy (anonymous):

Hey! So I'm confused by your question. Are the commas and numbers the answer choices?

OpenStudy (anonymous):

339+638=977 1801+73=1874 46.8+780=826.8 381-137=244 715-4812=-4097 930-38.7=891.3

OpenStudy (anonymous):

i am doing evaluate addition and subtraction expressions